OpenStudy (anonymous):

@Ray10 same matrix but A^-1 if you wouldnt mind :)

OpenStudy (ray10):

okay so A= \[\left[\begin{matrix}2 & 1 \\ 3 & 2\end{matrix}\right]\] correct? to work out the inverse you can follow this formula \[\frac{ 1 }{ ad-bc } (\left[\begin{matrix}d & -b \\ -c & a\end{matrix}\right]\]

OpenStudy (ray10):

to make it easier to understand the formula, with a 2x2 matrix, to find it's inverse, you switch the positions of "a" and "d" and times "b" and "c" by -1

OpenStudy (ray10):

can you work out the inverse from that? :)

OpenStudy (ray10):

@sky202 hope that helped!

OpenStudy (anonymous):

thanks :)